2. Write first four terms of the A.P. when the first term a and the common difference are given as follows:
(i) a = 10, d = 10
(ii) a = -2, d = 0
(iii) a = 4, d = – 3
(iv) a = -1 d = 1/2
(v) a = – 1.25, d = – 0.25

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) a = 10, d = 10 Let us consider, the Arithmetic Progression series be a1, a2, a3, a4, a5 … a1 = a = 10 a2 = a1+d = 10+10 = 20 a3 = a2+d = 20+10 = 30 a4 = a3+d = 30+10 = 40 a5 = a4+d = 40+10 = 50 And so on… Therefore, the A.P. series will be 10, 20, 30, 40, 50 … And First four terms of this A.P. will be 10, 20, 30, and 40. (ii) a = – 2, d = 0 Let us consider, the Arithmetic Progression series be a1, a2, a3, a4, a5 … a1 = a = -2 a2 = a1+d = – 2+0 = – 2 a3 = a2+d = – 2+0 = – 2 a4 = a3+d = – 2+0 = – 2 Therefore, the A.P. series will be – 2, – 2, – 2, – 2 … And, First four terms of this A.P. will be – 2, – 2, – 2 and – 2. (iii) a = 4, d = – 3 Let us consider, the Arithmetic Progression series be a1, a2, a3, a4, a5 … a1 = a = 4 a2 = a1+d = 4-3 = 1 a3 = a2+d = 1-3 = – 2 a4 = a3+d = -2-3 = – 5 Therefore, the A.P. series will be 4, 1, – 2 – 5 … And, first four terms of this A.P. will be 4, 1, – 2 and – 5. (iv) a = – 1, d = 1/2 Let us consider, the Arithmetic Progression series be a1, a2, a3, a4, a5 … a2 = a1+d = -1+1/2 = -1/2 a3 = a2+d = -1/2+1/2 = 0 a4 = a3+d = 0+1/2 = 1/2 Thus, the A.P. series will be-1, -1/2, 0, 1/2 And First four terms of this A.P. will be -1, -1/2, 0 and 1/2. (v) a = – 1.25, d = – 0.25 Let us consider, the Arithmetic Progression series be a1, a2, a3, a4, a5 … a1 = a = – 1.25 a2 = a1 + d = – 1.25-0.25 = – 1.50 a3 = a2 + d = – 1.50-0.25 = – 1.75 a4 = a3 + d = – 1.75-0.25 = – 2.00 Therefore, the A.P series will be 1.25, – 1.50, – 1.75, – 2.00 …….. And first four terms of this A.P. will be – 1.25, – 1.50, – 1.75 and – 2.00.